# The Absolute Anabelian Geometry of Virtual Curves Arising from Sections of Arithmetic Fundamental Groups of Configuration Spaces

### v1

@@ 1. Introduction
The objective of this paper is to introduce the notions of virtual varieties [cf. Definition 3.1] and pointed virtual curves [cf. Definition 3.5] (a special case of the notion of a virtual variety), and to study various anabelian reconstructions related to pointed virtual curves, especially, pointed virtual curves that arise from $2$-configuration spaces [cf. § 2, the discussion entitled “Configuration Spaces”]. Roughly speaking, given a family of pointed curves, the notion of a pointed virtual curve is a generalization of the notion of a fiber of that family over a rational point, where the rational point is replaced by a Galois section of the étale fundamental group of the base space of the family.
Under suitable assumptions, if the Galois section arises from a rational point $x$, then the virtual fundamental group associated to the pointed virtual curve is naturally isomorphic, up to composition with inner automorphisms, to the étale fundamental group of the fiber of $x$ [cf. Proposition 3.7]. Thus one may regard pointed curves as the special class of pointed virtual curves for which the given section is geometric. In particular, if one considers the family of pointed curves associated to a $2$-configuration space, then one may regard the pointed virtual curve that arises from a section as the “cuspidalization” of that section.
@@
[cf. Definition 3.1; § 2, the discussion entitled “Configuration Spaces”] that arises from the $2$-configuration space of a curve $X$ satisfying certain conditions and a given Galois section $s:G_{k}\to\Pi_{X}$.
@@
([cf. Theorem 8.2])
Let $k$ be either a number field or a mixed-characteristic local field; $\bar{k}$ an algebraic closure of $k$; $G_{k}$ the Galois group $\text{\rm Gal}(\bar{k}/k)$; $X$ a smooth, geometrically connected curve over $k$ of type $(0,r)$, where $r\geq 3$; $s:G_{k}\rightarrow\Pi_{X}$ a section of the natural surjection $\Pi_{X}\twoheadrightarrow G_{k}$. Write $\text{\rm pr}^{2/1}_{X}:X_{2}\rightarrow X$ for the first projection; $\mathbf{F}$ for the category of fields; $\Pi_{s}$ for the virtual fundamental group $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ [cf. Definition 3.1; § 2, the discussion entitled “Configuration Spaces”].
@@
from the abstract group $\Pi_{s}$ functorially with respect to isomorphisms of topological groups such that the action of $\Pi_{s}$ on $\text{\rm Base}(\Pi_{s})$ functorially induced by the conjugation action of $\Pi_{s}$ on itself factors through the quotient $\Pi_{s}\twoheadrightarrow G_{k}$. Moreover, there exists a $G_{k}$-equivariant isomorphism $\text{\rm Base}(\Pi_{s})\simeq\bar{k}$.
(ii) One may reconstruct an $\mathbf{F}$-graph [cf. § 2, the discussion entitled “Graphs and Categories”]
$\text{\rm Field}(\Pi_{s})=(\mathfrak{G}(\Pi_{s}),\mathbf{V}(\Pi_{s}),\mathbf{E}(\Pi_{s}))$
and a morphism of $\mathbf{F}$-graphs
$F\coloneqq(\text{\rm Id}_{\mathfrak{G}(\Pi_{s})},\tau):\text{\rm Base}(\Pi_{s})_{\mathfrak{G}(\Pi_{s})}\to\text{\rm Field}(\Pi_{s})$
(i.e., from the constant $\mathbf{F}$-graph associated to $\text{\rm Base}(\Pi_{s})$ over $\mathfrak{G}(\Pi_{s})$
$(\mathfrak{G}(\Pi_{s}),\mathbf{V}^{\prime}(\Pi_{s}),\mathbf{E}^{\prime}(\Pi_{s}))\coloneqq\text{\rm Base}(\Pi_{s})_{\mathfrak{G}(\Pi_{s})}$
to $\text{\rm Field}(\Pi_{s})$) from the abstract group $\Pi_{s}$ functorially with respect to isomorphisms of topological groups such that the following conditions are satisfied:
  1. (1)
    The action of $\Pi_{s}$ on $\text{\rm Field}(\Pi_{s})$ functorially induced by the conjugation action of $\Pi_{s}$ on itself factors through the quotient $\Pi_{s}\twoheadrightarrow G_{k}$. In particular, for each vertex $v\in V(\mathfrak{G}(\Pi_{s}))$, the stabilizer $\text{\rm Stab}(v)$ of $v$ in $G_{k}$ acts on $\mathbf{V}(\Pi_{s})(v)$.
  2. (2)
    For each vertex $v\in V(\mathfrak{G}(\Pi_{s}))$, there exists a $\text{\rm Stab}(v)$-equivariant commutative diagram
[…]
    where we write $K(X\times_{k}\bar{k})$ for the base change of the function field of $X$ to $\bar{k}$ and $f$ for the naturally induced inclusion of fields. Here the $\text{\rm Stab}(v)$-action on $K(X\times_{k}\bar{k})$ and $\bar{k}$ is induced by the Galois action of $G_{k}$.
(iii) The isomorphism class of the function field $K(X)$ is determined by the isomorphism class of the abstract profinite group $\Pi_{s}$.
(iv) Suppose that $r=3$. Recall that the group of automorphisms $\text{\rm Aut}(\Pi_{X})$ of the profinite group $\Pi_{X}$ acts naturally on the set of sections of the surjection $\Pi_{X}\twoheadrightarrow G_{k}$ [cf. [15], Theorem 2.6, (v), (vi)]. Then the $\text{\rm Aut}(\Pi_{X})$-orbit of the section $s$ is determined by the isomorphism class of the abstract profinite group $\Pi_{s}$. Moreover, the isomorphism class of the scheme $X$ is determined by the isomorphism class of the abstract profinite group $\Pi_{s}$.
@@
Let $k$ be a field of characteristic zero; $\mathcal{C}$ a pointed virtual curve [cf. Definition 3.5] of type $(g,r)$ with source $X$ and base $k$; $\bar{k}$ an algebraic closure of $k$; $G_{k}$ the Galois group $\text{\rm Gal}(\bar{k}/k)$. Assume that $r\geq 1$.
@@
(ii) Let $l$ be any prime such that $k$ is $l$-cyclotomically full [cf. Definition 4.10(ii)]. Consider the determinant map $\text{\rm Det}_{\Delta_{\mathcal{C}}}^{l}:\Pi_{\mathcal{C}}\rightarrow(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ [cf. Definition 4.10(iii)]. The homomorphism $(\text{\rm Det}_{\Delta_{\mathcal{C}}}^{l})^{\text{\rm ab-t}}:(\Pi_{\mathcal{C}})^{\text{\rm ab-t}}\rightarrow(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ induced on torsion-free quotients factors through $\Pi_{\mathcal{C}}\rightarrow G_{k}$, hence may be considered as a character of $G_{k}$. Then $w^{l}_{(\text{\rm cycl}_{G_{k}})^{\text{\rm ab-t}}}((\text{\rm Det}_{\Delta_{\mathcal{C}}})^{\text{\rm ab-t}})$ [cf. Definition 4.10(i)] exists and is independent of the section $s$.
(iii) Suppose that there exists at least one $l$ that satisfies the condition in (ii). Then one may reconstruct the set of conjugacy classes of cuspidal inertia subgroups $\text{\rm Cusp}(\mathcal{C})$ [cf. Definition 4.19] of $\Delta_{\mathcal{C}}$.
@@
[…]
@@
This paper is organized as follows: In § 3, introduce the notion of a virtual curve and discuss some basic properties of virtual curves. In § 4, we prove Theorem B by considering the Weil pairing associated to the Albanese variety [cf. [15], Appendix] of the curve under consideration and deduce a generalization of [15], Lemma 4.5. In § 5, we discuss various properties of configuration spaces and prove Theorem C by using a “stricter” family of open subgroups than the one in [13], Lemma 1.1.4. In § 6, we discuss how to “decuspidalize” the diagonal of configuration space so that we may reduce to the case of a fundamental group of a curve of strictly Belyi type and, by applying methods in [17], Chapter 1, reconstruct various copies of the function field of the curve, thus proving Theorem A(iii). In § 7, we explain how to reconstruct the section from the abstract virtual fundamental group, thus proving Theorem A(iv). In § 8, we organize various reconstructions in § 6 to formulate a functorial reconstruction of the $\mathbf{F}$-graph $\text{\rm Field}(\Pi_{[\text{\rm pr}^{2/1}_{X},s]})$, thus proving Theorem A(i), (ii).
@@ 2. Notations and Terminologies
Throughout the paper, we omit the basepoint since there exists a natural isomorphism between fundamental groups associated to different basepoints which is well-defined up to an inner automorphism. We denote the étale fundamental group of $X$ with basepoint omitted as $\Pi_{X}$. If $X$ is the spectrum of a field $k$, we also use the notation $G_{k}$ (i.e., the “absolute Galois group of $k$”) to denote $\Pi_{X}$. Let $X$ be a scheme that is geometrically connected over a field $k$. Then denote by $\Delta_{X}$ the kernel of the natural surjection $\Pi_{X}\twoheadrightarrow G_{k}$, i.e., the geometric fundamental group.
@@ 3. Virtual Curves
One approach to the cuspidalization problem is the method of configuration spaces. Roughly speaking, if one may reconstruct the morphism $(\text{\rm pr}^{2/1}_{X})_{*}:\Pi_{X_{2}}\to\Pi_{X}$ [cf. § 2, the discussion entitled “Configuration Spaces”] from $\Pi_{X}$ (where we note that this is in general not possible without additional assumptions), then one may reconstruct $\Pi_{X\setminus x}$ as the preimage of $s(G_{k})$ by $(\text{\rm pr}^{2/1}_{X})_{*}$, where $s:G_{k}\to\Pi_{X}$ is the section of the natural surjection $\Pi_{X}\twoheadrightarrow G_{k}$ that arises from the $k$-rational point $x$ [cf. Proposition 3.7].
@@
$1\rightarrow\Delta_{\mathcal{C}}\rightarrow\Pi_{X}\rightarrow\Pi_{Y}\rightarrow 1$
@@
${X^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}}$$\scriptstyle{h}$${Y^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${Y}$
@@
Note that the étale covering $\mathcal{C}^{\prime}\rightarrow\mathcal{C}$ of Corollary 3.4(ii) is not necessarily unique up to isomorphism.
@@
The following corollary follows immediately from Proposition 3.7.
@@ 4. Weights of Determinant Maps
In general, it is difficult to try to recover any sort of scheme structure from the virtual fundamental group. Nevertheless we note that if we work in the situation of Proposition 3.7, then the geometric virtual fundamental group is isomorphic to the geometric étale fundamental group of a certain curve, and it is natural to expect that it should be possible to reconstruct data corresponding to the cuspidal inertia subgroups [cf. Definition 4.19]. This makes sense since in the case where the section arises from a rational point [cf. Corollary 3.8], the set of conjugacy classes of cuspidal inertia subgroups, which arises from the set of conjugacy classes of decomposition groups determined by the cusps, is independent of the (geometric) section chosen [cf. Theorem 4.20]. Therefore, we may expect that even if we omit the assumption that the section arises from a rational point, it should still be possible to reconstruct data corresponding to the cuspidal inertia subgroups.
@@
Let $k$ be a field of characteristic zero; $\mathcal{G}$ a class of commutative group schemes of finite type over $k$. Suppose that $(X,x)$ is a pair where $X$ is a geometrically integral variety over $k$; $x$ is a $k$-rational point of $X$. Then we shall refer to a morphism $f:(X,x)\rightarrow(A,0_{A})$ of pointed $k$-varieties (i.e., “pointed variety over $k$” in the sense of [15], Definition A.1) as a $\mathcal{G}$-Albanese morphism if $A\in\mathcal{G}$, and for any morphism $f^{\prime}:(X,x)\rightarrow(A^{\prime},0_{A^{\prime}})$ of pointed $k$-varieties, where $A^{\prime}\in\mathcal{G}$, there exists a unique morphism of group schemes $g:A\rightarrow A^{\prime}$ such that $f^{\prime}=g\circ f$. We refer to $A$ as the $\mathcal{G}$-Albanese variety of $X$ (which in fact depends on the choice $x$ of a $k$-rational point of $X$). If $\mathcal{G}$ is the class of semi-abelian varieties [cf. [2], Chapter 1, Definition 2.3], we often abbreviate the term “$\mathcal{G}$-Albanese” as “Albanese”.
@@
The assertions of Proposition 4.2 follow immediately from [15], Proposition A.6, and [15], Proposition A.8, together with the fact that $H^{2}(X,\mathcal{O}_{X})=0$ (which implies that the Picard scheme of $X$ is smooth), and the fact that the abelianizations $(\Pi_{X_{\bar{k}}})^{\text{\rm ab}}$, $(\Pi_{Y_{\bar{k}}})^{\text{\rm ab}}$ of the geometric fundamental groups are torsion-free [cf. [18], Remark 1.2.2]. ∎
@@
[cf. Corollary 3.2] induces an (outer) isomorphism
@@
(ii) is a special case of Proposition 3.7. ∎
@@
(iv) Let $n$ be a positive integer. Then denote by $\mu_{n}(\bar{k})$ the group of roots of unity of order $n$ in $\bar{k}$. Note that $\mu_{n}(\bar{k})$ is naturally isomorphic to (${}_{n}(\mathbb{G}_{m,\bar{k}})$. Let $l$ be any prime number. Then denote by $\mu(\bar{k})$ (respectively, $\mu_{\mathbb{Z}_{l}}(\bar{k})$, $\mu_{\mathbb{Q}_{l}}(\bar{k})$) the Tate module $\text{\rm T}(\mathbb{G}_{m,\bar{k}})$ (respectively, $\text{\rm T}_{\mathbb{Z}_{l}}(\mathbb{G}_{m,\bar{k}})$, $\text{\rm T}_{\mathbb{Q}_{l}}(\mathbb{G}_{m,\bar{k}})$).
@@
of $\mathbb{Q}_{l}$-vector spaces with $G_{k}$-action. We refer to the pairing $\theta_{\mathbb{Z}_{l}}$ as the pro-${l}$ Weil pairing associated to $\theta$.
@@
Let $k$ be a field of characteristic zero; $\bar{k}$ an algebraic closure of $k$; $A$ a semi-abelian variety over $k$. Then there exists a natural isomorphism
@@
To show that $f$ is injective, it suffices to show that for any finite Galois étale covering $g:A^{\prime}\to A$, there exists an $n\in\mathbb{N}^{+}$ and an étale covering $h:A\to A^{\prime}$ such that $g\circ h=n_{A}$. By Lemma 4.8 below, there exists a commutative algebraic group structure on $A^{\prime}$ such that $g$ is a homomorphism of algebraic groups. Let $n$ be the degree of $g$. Then the kernel of $g$ is contained in the kernel of $n_{A^{\prime}}$. Therefore, there exists a morphism $h:A\to A^{\prime}$ such that $h\circ g=n_{A^{\prime}}$, so $g\circ h\circ g=g\circ n_{A^{\prime}}=n_{A}\circ g$. Since $g$ is dominant, this implies that $g\circ h=n_{A}$, as desired. ∎
@@
Note that it follows from the proof of Proposition 4.6 that, in the notation of this proof, $A^{\prime}$ is semi-abelian.
@@
[…]
where the vertical arrows are the natural homomorphisms. By Lemma 4.9 below, the horizontal sequences are exact. Since $T$ is a split torus, the left-hand vertical arrow is an isomorphism. Since $X$ is proper, the right-hand vertical arrow is an isomorphism [cf. [4], Exposé X, Corollaire 1.7]. Therefore, the middle vertical arrow determines a natural isomorphism $\Pi_{A\times_{k}A}\xrightarrow{\sim}\Pi_{A}\times\Pi_{A}$. A similar argument shows that we have a natural isomorphism $\Pi_{A\times_{k}A\times_{k}A}\xrightarrow{\sim}\Pi_{A}\times\Pi_{A}\times\Pi_{A}$.
@@
[…]
@@
[…]
@@
Finally, since for $n\in\mathbb{N}^{+}$, $n_{A}$ is a Galois étale covering [cf. [1], Exposé $\text{\rm VI}_{\text{\rm B}}$, Proposition 1.3, Corollaire 1.4.1] that restricts to $n_{T}$ on $T$, we conclude that the natural homomorphism $\Pi_{T}\to\Pi_{A}$ is injective. This completes the proof of Lemma 4.9. ∎
By Proposition 4.6, the étale fundamental group of a semi-abelian variety may be naturally identified with its Tate module. Next, we review the language of cyclotomic characters and weights.
@@
(i) Let $G$ be a profinite group; $l$ a prime number; $\chi_{G}$, $f$ two torsion-free characters (i.e., two continuous homomorphisms $G\rightarrow(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ of profinite groups) of $G$. Then we say that $f$ is of weight $w$ relative to $\chi_{G}$ if there exist $a,b\in\mathbb{Z}$, $b\neq 0$ , such that $w=\frac{2a}{b}$, and $(f^{\text{\rm ab-t}})^{b}=((\chi_{G})^{\text{\rm ab-t}})^{a}$. (Here, the “$2$” appears in order to ensure that the weights of the characters that we are interested in are integers. This convention is consistent with the usual conventions concerning weights that are applied in the context of the Riemann Hypothesis over finite fields.) In general, $f$ might not be of weight $w$ for any rational number $w$. On the other hand, if the image of $\chi_{G}$ is open, then there exists at most one rational number $w$, which we shall denote $w_{\chi_{G}}(f)$ when it exists, such that $f$ is of weight $w$ relative to $\chi_{G}$.

### v2

@@ 1. Introduction
The objective of this paper is to introduce the notions of virtual varieties [cf. Definition˜3.1] and pointed virtual curves [cf. Definition˜3.5] (a special case of the notion of a virtual variety), and to study various anabelian reconstructions related to pointed virtual curves, especially, pointed virtual curves that arise from $2$-configuration spaces [cf. §˜2, the discussion entitled “Configuration Spaces”]. Roughly speaking, given a family of pointed curves, the notion of a pointed virtual curve is a generalization of the notion of a fiber of that family over a rational point, where the rational point is replaced by a Galois section of the étale fundamental group of the base space of the family.
Under suitable assumptions, if the Galois section arises from a rational point $x$, then the virtual fundamental group associated to the pointed virtual curve is naturally isomorphic, up to composition with inner automorphisms, to the étale fundamental group of the fiber of $x$ [cf. Proposition˜3.7]. Thus one may regard pointed curves as the special class of pointed virtual curves for which the given section is geometric. In particular, if one considers the family of pointed curves associated to a $2$-configuration space, then one may regard the pointed virtual curve that arises from a section as the “cuspidalization” of that section.
@@
[cf. Definition˜3.1; §˜2, the discussion entitled “Configuration Spaces”] that arises from the $2$-configuration space of a curve $X$ satisfying certain conditions and a given Galois section $s:G_{k}\to\Pi_{X}$.
@@
([cf. Theorem 6.14(iii); Theorem 6.15(vi)(v)(vi)])
Let $k$ be either a number field or a mixed-characteristic local field; $\bar{k}$ an algebraic closure of $k$; $G_{k}$ the Galois group $\text{\rm Gal}(\bar{k}/k)$; $X$ a smooth, geometrically connected curve over $k$ of type $(0,r)$, where $r\geq 3$; $s:G_{k}\rightarrow\Pi_{X}$ a section of the natural surjection $\Pi_{X}\twoheadrightarrow G_{k}$. Write $\text{\rm pr}^{2/1}_{X}:X_{2}\rightarrow X$ for the first projection; $\mathbf{F}$ for the category of fields; $\Pi_{s}$ for the virtual fundamental group $\Pi_{[\text{\rm pr}^{2/1}_{X},s]}$ [cf. Definition˜3.1; §˜2, the discussion entitled “Configuration Spaces”].
@@
from the abstract profinite group $\Pi_{s}$ functorially with respect to isomorphisms of topological groups such that the action of $\Pi_{s}$ on $\text{\rm Base}(\Pi_{s})$ functorially induced by the conjugation action of $\Pi$ on itself factors through the quotient $\Pi_{s}\twoheadrightarrow G_{k}$. Moreover, there exists a $G_{k}$-equivariant isomorphism $\text{\rm Base}(\Pi_{s})\xrightarrow{\sim}\bar{k}$.
(ii) The isomorphism class of the function field $K(X)$ is determined by the isomorphism class of the abstract profinite group $\Pi_{s}$.
(iii) Suppose that $r=3$. Then the isomorphism class of the scheme $X$ is determined by the isomorphism class of the abstract profinite group $\Pi_{s}$.
@@
Let $k$ be a field of characteristic zero; $\mathcal{C}$ a pointed virtual curve [cf. Definition˜3.5] of type $(g,r)$ with source $X$ and base $k$; $\bar{k}$ an algebraic closure of $k$; $G_{k}$ the Galois group $\text{\rm Gal}(\bar{k}/k)$. Assume that $r\geq 1$.
@@
(ii) Let $l$ be any prime such that $k$ is $l$-cyclotomically full [cf. Definition˜4.10(ii)]. Consider the determinant map $\text{\rm Det}_{\Delta_{\mathcal{C}}}^{l}:\Pi_{\mathcal{C}}\rightarrow(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ [cf. Definition˜4.10(iii)]. The homomorphism $(\text{\rm Det}_{\Delta_{\mathcal{C}}}^{l})^{\text{\rm ab-t}}:(\Pi_{\mathcal{C}})^{\text{\rm ab-t}}\rightarrow(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ induced on torsion-free quotients factors through $\Pi_{\mathcal{C}}\rightarrow G_{k}$, hence may be considered as a character of $G_{k}$. Then $w^{l}_{(\text{\rm cycl}_{G_{k}})^{\text{\rm ab-t}}}((\text{\rm Det}_{\Delta_{\mathcal{C}}})^{\text{\rm ab-t}})$ [cf. Definition˜4.10(i)] exists and is independent of the section $s$.
(iii) Suppose that there exists at least one $l$ that satisfies the condition in (ii). Then one may reconstruct the set of conjugacy classes of cuspidal inertia subgroups $\text{\rm Cusp}(\mathcal{C})$ [cf. Definition˜4.19] of $\Delta_{\mathcal{C}}$.
@@
$\text{``semi-absolute-anabelian''}\Rightarrow\text{``absolute-anabelian''}$
@@
This paper is organized as follows: In §˜3, introduce the notion of a virtual curve and discuss some basic properties of virtual curves. In §˜4, we prove Theorem B by considering the Weil pairing associated to the Albanese variety [cf. [14], Appendix] of the curve under consideration and deduce a generalization of [14], Lemma 4.5. In §˜5, we discuss various properties of configuration spaces and prove Theorem C by using a “stricter” family of open subgroups than the one in [12], Lemma 1.1.4. In §˜6, we discuss how to “decuspidalize” the diagonal of configuration space so that we may reduce to the case of a fundamental group of a curve of strictly Belyi type and, by applying methods in [16], Chapter 1, reconstruct various copies of the function field of the curve, thus proving Theorem A.
@@ 2. Notations and Terminologies
Throughout the paper, we omit the basepoint since there exists a natural isomorphism between fundamental groups associated to different basepoints which is well-defined up to an inner automorphism. We denote the étale fundamental group of $X$ with basepoint omitted as $\Pi_{X}$. If $X$ is the spectrum of a field $k$, we also use the notation $G_{k}$ (i.e., the “absolute Galois group of kk”) to denote $\Pi_{X}$. Let $X$ be a scheme that is geometrically connected over a field $k$. Then denote by $\Delta_{X}$ the kernel of the natural surjection $\Pi_{X}\twoheadrightarrow G_{k}$, i.e., the geometric fundamental group.
@@ 3. Virtual Curves
One approach to the cuspidalization problem is the method of configuration spaces. Roughly speaking, if one may reconstruct the morphism $(\text{\rm pr}^{2/1}_{X})_{*}:\Pi_{X_{2}}\to\Pi_{X}$ [cf. §˜2, the discussion entitled “Configuration Spaces”] from $\Pi_{X}$ (where we note that this is in general not possible without additional assumptions), then one may reconstruct $\Pi_{X\setminus x}$ as the preimage of $s(G_{k})$ by $(\text{\rm pr}^{2/1}_{X})_{*}$, where $s:G_{k}\to\Pi_{X}$ is the section of the natural surjection $\Pi_{X}\twoheadrightarrow G_{k}$ that arises from the $k$-rational point $x$ [cf. Proposition˜3.7].
@@
$1\rightarrow\Delta_{\mathcal{C}}\rightarrow\Pi_{X}\rightarrow\Pi_{Y}\rightarrow 1.$
@@
[…]
@@
Note that the étale covering $\mathcal{C}^{\prime}\rightarrow\mathcal{C}$ of Corollary˜3.4(ii) is not necessarily unique up to isomorphism.
@@
The following corollary follows immediately from Proposition˜3.7.
@@ 4. Weights of Determinant Maps
In general, it is difficult to try to recover any sort of scheme structure from the virtual fundamental group. Nevertheless we note that if we work in the situation of Proposition˜3.7, then the geometric virtual fundamental group is isomorphic to the geometric étale fundamental group of a certain curve, and it is natural to expect that it should be possible to reconstruct data corresponding to the cuspidal inertia subgroups [cf. Definition˜4.19]. This makes sense since in the case where the section arises from a rational point [cf. Corollary˜3.8], the set of conjugacy classes of cuspidal inertia subgroups, which arises from the set of conjugacy classes of decomposition groups determined by the cusps, is independent of the (geometric) section chosen [cf. Theorem˜4.20]. Therefore, we may expect that even if we omit the assumption that the section arises from a rational point, it should still be possible to reconstruct data corresponding to the cuspidal inertia subgroups.
@@
Let $k$ be a field of characteristic zero; $\mathcal{G}$ a class of commutative group schemes of finite type over $k$. Suppose that $(X,x)$ is a pair where $X$ is a geometrically integral variety over $k$; $x$ is a $k$-rational point of $X$. Then we shall refer to a morphism $f:(X,x)\rightarrow(A,0_{A})$ of pointed $k$-varieties (i.e., “pointed variety over kk” in the sense of [14], Definition A.1) as a $\mathcal{G}$-Albanese morphism if $A\in\mathcal{G}$, and for any morphism $f^{\prime}:(X,x)\rightarrow(A^{\prime},0_{A^{\prime}})$ of pointed $k$-varieties, where $A^{\prime}\in\mathcal{G}$, there exists a unique morphism of group schemes $g:A\rightarrow A^{\prime}$ such that $f^{\prime}=g\circ f$. We refer to $A$ as the $\mathcal{G}$-Albanese variety of $X$ (which in fact depends on the choice $x$ of a $k$-rational point of $X$). If $\mathcal{G}$ is the class of semi-abelian varieties [cf. [2], Chapter 1, Definition 2.3], we often abbreviate the term $\mathcal{G}$-Albanese” as “Albanese”.
@@
The assertions of Proposition˜4.2 follow immediately from [14], Proposition A.6, and [14], Proposition A.8, together with the fact that $H^{2}(X,\mathcal{O}_{X})=0$ (which implies that the Picard scheme of $X$ is smooth), and the fact that the abelianizations $(\Pi_{X_{\bar{k}}})^{\text{\rm ab}}$, $(\Pi_{Y_{\bar{k}}})^{\text{\rm ab}}$ of the geometric fundamental groups are torsion-free [cf. [17], Remark 1.2.2]. ∎
@@
[cf. Corollary˜3.2] induces an (outer) isomorphism
@@
(ii) is a special case of Proposition˜3.7. ∎
@@
[…]
@@
of $\mathbb{Q}_{l}$-vector spaces with $G_{k}$-action. We refer to the pairing $\theta_{\mathbb{Z}_{l}}$ as the pro-l{l} Weil pairing associated to $\theta$.
@@
Let $k$ be a field of characteristic zero; $\bar{k}$ an algebraic closure of $k$; AA a semi-abelian variety over $k$. Then there exists a natural isomorphism
@@
To show that $f$ is injective, it suffices to show that for any finite Galois étale covering $g:A^{\prime}\to A$, there exists an $n\in\mathbb{N}^{+}$ and an étale covering $h:A\to A^{\prime}$ such that $g\circ h=n_{A}$. By Lemma˜4.8 below, there exists a commutative algebraic group structure on $A^{\prime}$ such that $g$ is a homomorphism of algebraic groups. Let $n$ be the degree of $g$. Then the kernel of $g$ is contained in the kernel of $n_{A^{\prime}}$. Therefore, there exists a morphism $h:A\to A^{\prime}$ such that $h\circ g=n_{A^{\prime}}$, so $g\circ h\circ g=g\circ n_{A^{\prime}}=n_{A}\circ g$. Since $g$ is dominant, this implies that $g\circ h=n_{A}$, as desired. ∎
@@
Note that it follows from the proof of Proposition˜4.6 that, in the notation of this proof, $A^{\prime}$ is semi-abelian.
@@
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Pi_{T\times_{k}T}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Pi_{A\times_{k}A}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Pi_{X\times_{k}X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Pi_{T}\times\Pi_{T}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Pi_{A}\times\Pi_{A}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Pi_{X}\times\Pi_{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1,}$
where the vertical arrows are the natural homomorphisms. By Lemma˜4.9 below, the horizontal sequences are exact. Since $T$ is a split torus, the left-hand vertical arrow is an isomorphism. Since $X$ is proper, the right-hand vertical arrow is an isomorphism [cf. [4], Exposé X, Corollaire 1.7]. Therefore, the middle vertical arrow determines a natural isomorphism $\Pi_{A\times_{k}A}\xrightarrow{\sim}\Pi_{A}\times\Pi_{A}$. A similar argument shows that we have a natural isomorphism $\Pi_{A\times_{k}A\times_{k}A}\xrightarrow{\sim}\Pi_{A}\times\Pi_{A}\times\Pi_{A}$.
@@
${B\times_{k}B\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{B}}$$\scriptstyle{f^{B}_{A}\times_{k}f^{B}_{A}}$${B\times_{k}B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{B}_{A}\times_{k}f^{B}_{A}}$${A\times_{k}A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{A}}$${A\times_{k}A}$
@@
${B\times_{k}B\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h_{1}}$$\scriptstyle{h_{2}}$${B\times_{k}B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{B}_{A}\times_{k}f^{B}_{A}}$${A\times_{k}A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{A}}$${A\times_{k}A}$
@@
Finally, since for $n\in\mathbb{N}^{+}$, $n_{A}$ is a Galois étale covering [cf. [1], Exposé $\text{\rm VI}_{\text{\rm B}}$, Proposition 1.3, Corollaire 1.4.1] that restricts to $n_{T}$ on $T$, we conclude that the natural homomorphism $\Pi_{T}\to\Pi_{A}$ is injective. This completes the proof of Lemma˜4.9. ∎
By Proposition˜4.6, the étale fundamental group of a semi-abelian variety may be naturally identified with its Tate module. Next, we review the language of cyclotomic characters and weights.
@@
(i) Let $G$ be a profinite group; $l$ a prime number; $\chi_{G}$, $f$ two torsion-free characters (i.e., two continuous homomorphisms $G\rightarrow(\mathbb{Z}_{l}^{\times})^{\text{\rm tf}}$ of profinite groups) of $G$. Then we say that $f$ is of weight $w$ relative to $\chi_{G}$ if there exist $a,b\in\mathbb{Z}$, $b\neq 0$ , such that $w=\frac{2a}{b}$, and $(f^{\text{\rm ab-t}})^{b}=((\chi_{G})^{\text{\rm ab-t}})^{a}$. (Here, the “22” appears in order to ensure that the weights of the characters that we are interested in are integers. This convention is consistent with the usual conventions concerning weights that are applied in the context of the Riemann Hypothesis over finite fields.) In general, $f$ might not be of weight $w$ for any rational number $w$. On the other hand, if the image of $\chi_{G}$ is open, then there exists at most one rational number $w$, which we shall denote $w_{\chi_{G}}(f)$ when it exists, such that $f$ is of weight $w$ relative to $\chi_{G}$.